# Orthogonal projections of hypercubes

### v3

@@ VII Applications
VII.1 The finite artificial kagome spin-ice system
@@
The Hamiltonian of the Ising spin system is defined as $\mathcal{H}\left(\bm{s}\right)\coloneqq-\frac{1}{2}\sum_{i=1}^{N}\sum_{j=1}^{N}s_{i}J_{i,j}s_{j}-\sum_{i=1}^{N}s_{i}h_{i}=-\frac{1}{2}\bm{s}^{\top}\bm{J}\bm{s}-\bm{s}^{\top}\bm{h},$ where the state vector is defined as $\bm{s}\coloneqq\begin{bmatrix}s_{1}&\cdots&s_{N}\end{bmatrix}^{\top}$ with the element as the Ising spin variable (up, $\uparrow$, or down, $\downarrow$), $s_{i}\in\left\{\bullet=~{}\uparrow~{}\coloneqq+1,\circ=~{}\downarrow~{}\coloneqq-1\right\}$. A state $\bm{s}$ is a vertex of the $N$-dimensional hypercube, and a state transition with a single spin flip corresponds to hypercubic edge. The element of the interaction matrix $\bm{J}\in\left\{-1,0\right\}^{N\times N}$ is $J_{i,j}=J_{j,i}=-1$ when spin $i$ and $j$ antiferromagnetically interact (antialignment of spin $i$ and $j$ decreases energy) on the kagome lattice of Fig. 9(b) and $J_{i,j}=J_{j,i}=0$ otherwise. Self-interactions do not exist ($J_{i,i}=0$, $\forall i$). The external magnetic field $\bm{h}\coloneqq\begin{bmatrix}h_{1}&\cdots&h_{N}\end{bmatrix}^{\top}\in\mathbb{R}^{N}$ is assigned to be zero $\bm{0}$ in this study.
@@
We examine the spin-ice systems through the interaction network and associated frustration. The interaction network of the one-ring system [Fig. 9(b1)] is a ring without frustration, $\varPhi\left(\mathcal{C}\right)\nless 0$, and there are two ground states: a ground state $\begin{bmatrix}\circ&\bullet&\bullet&\circ&\circ&\bullet\end{bmatrix}^{\top}$ and its global spin-flipped state $\begin{bmatrix}\bullet&\circ&\circ&\bullet&\bullet&\circ\end{bmatrix}^{\top}$. The two-ring system has several ground states owing to the frustrated interaction involving spin 6, such as $\mathcal{C}=\left\{J_{6,10},J_{10,9},J_{9,6}\right\}$ and $\mathcal{C}=\left\{J_{6,3},J_{3,2},J_{2,6}\right\}$. In the three-ring system, there are more frustrated cycles containing spins 6, 9, and 10, e.g., $\mathcal{C}=\left\{J_{6,10},J_{10,9},J_{9,6}\right\}$, $\mathcal{C}=\left\{J_{6,3},J_{3,2},J_{2,6}\right\}$, $\mathcal{C}=\left\{J_{9,8},J_{8,12},J_{12,9}\right\}$, and $\mathcal{C}=\left\{J_{10,13},J_{13,11},J_{11,10}\right\}$.
VII.2 Projecting the hypercubic energy landscape of the kagome spin-ice system
@@
For the projection of the hypercubic energy landscape, we calculate the covariance matrix [Eq. (7)] with the probability distribution of each state. The probability distribution of each state (hypercubic vertex) is determined by the canonical ensemble: $p\left(\bm{s}\right)=\frac{1}{Z}\exp\left[-\beta\mathcal{H}\left(\bm{s}\right)\right]$, where $Z\coloneqq\sum_{\bm{s}}\exp\left[-\beta\mathcal{H}\left(\bm{s}\right)\right]$ is the partition function. The temperature $T$ of the reservoir (bath) or inverse temperature $\beta$ is assigned to be $k_{\mathrm{B}}T=\frac{1}{\beta}=0.3$ in our projection. Here, $k_{\mathrm{B}}$ is the Boltzmann constant. Note that the probability distribution is symmetric with respect to the origin, $p\left(-\bm{s}\right)=p\left(\bm{s}\right)$ when the external field is zero, $\bm{h}=\bm{0}$, because the Hamiltonian has symmetry under the global spin flip, $\bm{s}\leftarrow-\bm{s}$, or $\mathbb{Z}_{2}$ symmetry, $\mathcal{H}\left(-\bm{s}\right)=-\frac{1}{2}\left(-\bm{s}\right)^{\top}\bm{J}\left(-\bm{s}\right)=-\frac{1}{2}\bm{s}^{\top}\bm{J}\bm{s}=\mathcal{H}\left(\bm{s}\right)$. This symmetry of the probability distribution makes the mean state vector the zero vector, $\bm{s}\right>=\bm{0}$ because $\bm{s}\right>=\sum_{\bm{s}}p\left(\bm{s}\right)\bm{s}=\sum_{-\bm{s}}p\left(-\bm{s}\right)\left(-\bm{s}\right)=\sum_{\bm{s}}p\left(\bm{s}\right)\left(-\bm{s}\right)=-\left<\bm{s}\right>$. Thus, the covariance matrix becomes $\bm{\varSigma}=\sum_{\bm{s}}p\left(\bm{s}\right)\bm{s}\bm{s}^{\top}$. The non-diagonal element (covariance) is $\varSigma_{i,j}=\sum_{\bm{s}}p\left(\bm{s}\right)s_{i}s_{j}$, and the diagonal element (variance) is $\varSigma_{i,i}=\sum_{\bm{s}}p\left(\bm{s}\right)s_{i}^{2}=\sum_{\bm{s}}p\left(\bm{s}\right)=1$. Therefore, the correlation between spin $i$ and $j$ is the covariance between them, $\frac{\varSigma_{i,j}}{\sqrt{\varSigma_{i,i}}\sqrt{\varSigma_{j,j}}}=\varSigma_{i,j}$. The covariance matrix becomes the correlation matrix. PCA finds the most correlated direction in the hypercubic state space.
In Figs. 10(a1), 10(a2), and 10(a3), the hypercubic energy landscapes of finite artificial kagome Ising spin systems are projected using PCA. Each state $\bm{s}$ corresponds to a vertex of the hypercube, and a state transition with a single spin flip corresponds to the edge of the hypercube. Thus, a hypercubic energy landscape is a kind of state transition diagram. For clarity of visualization, we show only the vertices with the lowest and the second-lowest energy, i.e., the ground states and the first-excited states, out of $2^{N}$ vertices (states) in Figs. 10(a1), 10(a2), and 10(a3). By visualizing hypercubic edges connecting the ground states and first-excited states, one can observe the pathways expected to be followed in the dynamics at low temperatures.
@@
Figure 10: Hypercubic energy landscape of finite artificial spin-ice systems and interaction networks colored by the angle of the biplot basis. (a) The hypercubic energy landscape of one-ring (a1), two-ring (a2), and three-ring (a3) systems by PC1 and PC2. For clarity of the figure, only the states with the lowest and the second-lowest energy are shown. The energy of the state colors vertices (blue for the lowest energy and red for the second-lowest energy). Biplot basis are shown on the top right, and their color indicates the angle of the biplot basis. Notice that in (a2) and (a3), several biplot basis are overlapped. One example pathway from a ground state to another ground state is shown as colored hypercubic edges in (a1). The colors of these lines are the same as the color of the corresponding biplot basis. Examples of states are shown as insets in (a1), (a2), and (a3). See the caption of Fig. 9 for the definition of state visualization on the interaction network. (b) PC1 loading of one-ring (b1), two-ring (b2), and three-ring (b3) systems. The color of the bar matches the color of the corresponding biplot basis in (a). (c) Same as (b) but for PC2. (d) The Ising spin interaction networks of one-ring (d1), two-ring (d2), and three-ring (d3) systems, where the nodes are colored by the angle of the corresponding biplot basis in (a).
VII.3 The hypercubic energy landscape of the one-ring system
@@
We first examine the one-ring system of Figs. 9(a1) and 9(b1). In Fig. 10(a1), we show the possible pathway from a ground state $\begin{bmatrix}\circ&\bullet&\bullet&\circ&\circ&\bullet\end{bmatrix}^{\top}$ to the globally spin-flipped state $\begin{bmatrix}\bullet&\circ&\circ&\bullet&\bullet&\circ\end{bmatrix}^{\top}$ by connecting the vertices with the lowest energy and the second-lowest energy using hypercubic edges. (See Figs. 9(b1) or 10(d1) for labeling of spins.) In this example, the hypercubic edge corresponds to the state transition with a single spin flip. The biplot basis are shown at the top right of Fig. 10(a1), and its color indicates the angle. We emphasize one particular pathway between two ground states by coloring the edges following the biplot basis in Fig. 10(b1). We found the state transition pathway is collective: the flipping order is restricted to keep the energy low. In a one-ring system, there are $6!=720$ possible pathways from one ground state to another ground state, but the number of most probable pathways is limited, as shown in Fig. 10(a1), and it is reduced to $966\times 2^{4}\times 1=$ 96 [70]. This reduction arises from the interaction network of the system and the associated hypercubic energy landscape.
@@
In addition, one can infer which spin is involved in which state transition by the biplot basis. One can know which spin flips in the state transition by comparing the length and angle of the hypercubic edge of interest with that of the biplot basis. The direction of the arrow of the biplot basis indicates the direction of the spin flip. In other words, if the state transition is along the direction of the arrow of the biplot basis, the spin flips from down to up. Otherwise, the spin flips from up to down. Notice that Fig. 10(a1) is the Hamming projection because the probability distribution is roughly the same as Eq. (11).
@@
To interpret the relation between PC loading and the interaction network, we visualize the loading of PC1 and PC2 in Figs. 10(b1) and 10(c1), respectively, colored by the angle of the biplot basis in Fig. 10(a1). We confirmed that the element-wise sign of PC1 corresponds to the ground state [Fig. 10(b1)]. The element-wise sign of PC2 loading corresponds to the state perpendicular to the ground state, which has a Hamming distance of $\frac{N}{2}$ from the ground state [Fig. 10(c1)].
@@
The interaction network in Fig. 10(d1), where nodes are colored by the angle of the biplot basis, shows the correlation captured by the first two PC loadings. Spins 1 and 5 have the same color, and spin 3, interacting with both spins 1 and 5, has the color complementary to that of spins 1 and 5. We can argue the same for spins 2, 4, and 6. Hence, the resulting hypercubic energy landscape is drawn to emphasize the correlation arising from the interaction.
VII.4 The hypercubic energy landscape of the two-ring system
@@
The hypercubic energy landscape of the two-ring system, which has several ground states, is illustrated in Fig. 10(a2). We show six ground states and pathways between them in a two-ring system with a biplot basis on the top right. Similar to the one-ring system in Fig. 10(a1), the most probable pathways between the ground states are shown by connecting the vertices with the lowest and the second-lowest energy. We discuss two pathways connecting the ground states as examples.
@@
The second example is the transition pathway from the ground state on the upper left side (around $\begin{bmatrix}-3&0.3\end{bmatrix}^{\top}$), to the ground state on the top (around $\begin{bmatrix}0&3.2\end{bmatrix}^{\top}$), i.e., state transition from $\begin{bmatrix}\circ&\bullet&\circ&\bullet&\bullet&\bullet&\circ&\circ&\bullet&\circ&\bullet\end{bmatrix}^{\top}$ to $\begin{bmatrix}\bullet&\circ&\circ&\bullet&\circ&\bullet&\circ&\bullet&\circ&\circ&\bullet\end{bmatrix}^{\top}$, which requires the flipping of spins 1, 2, 5, 8, and 9 [see biplot basis of Fig. 10(a2) and color of Figs. 10(b2), 10(c2), and 10(d2)]. This projection of the hypercube indicates that state transition happens with the correlated spin flips on the left-half part of the system (spins 1, 2, 5, 8, and 9). Among the left-half part of the system, spins 2 and 9 are more likely to flip at the beginning of the pathway. This difference arises from the unsatisfied interactions: $J_{3,9}$ and $J_{6,2}$. Spins 2 and 9 receive weaker local fields than spins 1, 5, and 8 because they have interactions involving spins 3, 6, and 10. Thus, spins 2 and 9 are expected to flip more than spins 1, 5, and 8. Because of the frustrated interactions involving spins 3, 6, and 10, spins 2 or 9 are the most probable to flip first. Then, the flips happen in the order of spins 2, 1, 5, 8, and 9 if spin 2 flips first, or in the order of spins 9, 8, 5, 1, and 2 if spin 9 flips first. Similar to the one-ring system in Fig. 10(a1), the number of most probable pathways is limited, but the constraints are stricter and fewer pathways are possible.
@@
To deepen our understanding of PC loading and interaction networks, we show the loading of PC1 and PC2 in Figs. 10(b2) and 10(c2), coloring them by the angle of the biplot basis in Fig. 10(a2). Again, we confirmed that the sign of PC1 loading [Fig. 10(b2)] corresponds to the ground state with the largest magnitude of the PC1 score, but spin 6 is not determined. Thus, the frustrated dynamics of spin 6 are captured by PC1. The element-wise sign of PC2 loading [Fig. 10(c2)] corresponds to the ground state with the largest magnitude of the PC2 score. Notice that the sixth element of PC2 loading has a slightly higher magnitude than the others, indicating the importance of spin 6. If spin 6 flips when the system is in the ground state $\operatorname{sgn}\left(\bm{u}_{2}\right)$, the energy of the system increases more than when another spin flips.
@@
The interaction network with nodes colored by the angle of the biplot basis in Fig. 10(d2) shows the inter-spin correlation captured by the biplot basis. The biplot basis capture the (anti)correlation of spins 1, 2, 5, 8, and 9 [yellow and blue arrows in Fig. 10(a2) top right] which corresponds to the left ring of the interaction network in Fig. 10(d2). The biplot basis also capture the (anti)correlation of spins 3, 4, 7, 10, and 11 [green and purple arrows in Fig. 10(a2) top right] which corresponds to the right ring of the interaction network in Fig. 10(d2). The hypercubic energy landscape using PCA provides insight into the dynamics through the correlation of spins.
VII.5 The hypercubic energy landscape of the three-ring system

### v4

@@ VII Applications
VII.1 Related studies and background
Several previous works have applied PCA to models in statistical mechanics, particularly in the study of phase transitions. Studies [100, 101, 102] demonstrate that PCA can successfully identify the order parameter of spin systems from sampled states. Moreover, PCA has been shown to detect order parameters in off-lattice [118] and active matter [119] systems, highlighting its applicability across a broad range of many-body systems. Thus, applying PCA enables the discovery of a system’s order parameter—even when the underlying Hamiltonian is unknown—which is valuable for the unsupervised detection of phase transitions.
Similar to previous studies [100, 101, 102], in this Sec. VII, we perform PCA on the states of Ising spin systems in a canonical ensemble. Our aim is, however, not only to visualize the states (hypercubic vertices) but also to capture state transitions (hypercubic edges). We seek to provide an alternative interpretation of PCA, focusing on state transition dynamics through correlated spin flips, rather than merely extracting meaningful parameters that describe the phase of the system. We begin by visualizing the hypercubic energy landscapes of Ising spin systems, which offer insights into the expected state transition dynamics at low temperatures—dynamics that are experimentally observable [52]. Using these hypercubic energy landscapes, we then directly investigate state transition dynamics by introducing the time dependence of the probability distribution, revealing state transition pathways as constrained probability fluxes that emerge from correlated spin flips.
@@ VII.1 Related studies and background
VII.2 The finite artificial kagome spin-ice system
@@
The Hamiltonian of the Ising spin system is defined as
$\mathcal{H}\left(\bm{s}\right)$ $\coloneqq-\frac{1}{2}\sum_{i=1}^{N}\sum_{j=1}^{N}s_{i}J_{i,j}s_{j}-\sum_{i=1}^{N}s_{i}h_{i}$ (68)
$=-\frac{1}{2}\bm{s}^{\top}\bm{J}\bm{s}-\bm{s}^{\top}\bm{h},$ (69)
where the state vector is defined as $\bm{s}\coloneqq\begin{bmatrix}s_{1}&\cdots&s_{N}\end{bmatrix}^{\top}$ with the element as the Ising spin variable (up $\uparrow$ or down $\downarrow$), $s_{i}\in\left\{\bullet=\,\uparrow\,\coloneqq+1,\circ=\,\downarrow\,\coloneqq-1\right\}$. A state $\bm{s}$ is a vertex of the $N$-dimensional hypercube, and a state transition with a single spin flip corresponds to hypercubic edge. The element of the interaction matrix $\bm{J}\in\left\{-1,0\right\}^{N\times N}$ is $J_{i,j}=J_{j,i}=-1$ when spin $i$ and $j$ antiferromagnetically interact (antialignment of spin $i$ and $j$ decreases energy) on the kagome lattice of Figs. 9(b1)–9(b3), and $J_{i,j}=J_{j,i}=0$ otherwise. Self-interactions do not exist ($J_{i,i}=0$, $\forall i$). The external magnetic field $\bm{h}\coloneqq\begin{bmatrix}h_{1}&\cdots&h_{N}\end{bmatrix}^{\top}\in\mathbb{R}^{N}$ is assigned to be zero $\bm{0}$ for projection of the hypercubic energy landscape, but later we will consider the case with a non-zero field.
@@
We examine the spin-ice systems through the interaction network and associated frustration. The interaction network of the one-ring system [Fig. 9(b1)] is a ring without frustration, $\varPhi\left(\mathcal{C}\right)\nless 0$, and there are two ground states: a ground state $\begin{bmatrix}\circ&\bullet&\bullet&\circ&\circ&\bullet\end{bmatrix}^{\top}$ and its global spin-flipped state $\begin{bmatrix}\bullet&\circ&\circ&\bullet&\bullet&\circ\end{bmatrix}^{\top}$. The two-ring system has several ground states owing to the frustrated interaction involving spin 6, such as $\mathcal{C}=\left\{J_{6,10},J_{10,9},J_{9,6}\right\}$ and $\mathcal{C}=\left\{J_{6,3},J_{3,2},J_{2,6}\right\}$. In the three-ring system, there are more frustrated cycles containing spins 6, 9, and 10, e.g., $\mathcal{C}=\left\{J_{6,10},J_{10,9},J_{9,6}\right\}$, $\mathcal{C}=\left\{J_{6,3},J_{3,2},J_{2,6}\right\}$, $\mathcal{C}=\left\{J_{9,8},J_{8,12},J_{12,9}\right\}$, and $\mathcal{C}=\left\{J_{10,13},J_{13,11},J_{11,10}\right\}$. Thus, the one-ring system has an unfrustrated interaction network, and the two- and three-ring systems have frustrated interaction networks.
VII.3 Projecting the hypercubic energy landscape of the kagome spin-ice system
For the projection of the hypercubic energy landscape, we calculate the covariance matrix [Eq. (9)] with the probability distribution of each state. The probability distribution of each state (hypercubic vertex) is determined by the canonical ensemble: Boltzmann distribution
$p\left(\bm{s}\right)=\frac{1}{\mathcal{Z}}\exp\left[-\beta\mathcal{H}\left(\bm{s}\right)\right],$ (70)
@@ VII.3 Projecting the hypercubic energy landscape of the kagome spin-ice system
where $\mathcal{Z}\coloneqq\sum_{\bm{s}}\exp\left[-\beta\mathcal{H}\left(\bm{s}\right)\right]$ is the partition function. The temperature $T$ of the reservoir (bath) or inverse temperature $\beta$ is assigned to be $k_{\mathrm{B}}T=\frac{1}{\beta}=0.3$ in our projection. Here, $k_{\mathrm{B}}$ is the Boltzmann constant. Note that the probability distribution is an even function, $p\left(-\bm{s}\right)=p\left(\bm{s}\right)$ when the external field is zero, $\bm{h}=\bm{0}$, because the Hamiltonian has symmetry under the global spin flip, $\bm{s}\leftarrow-\bm{s}$, or $\mathbb{Z}_{2}$ symmetry, $\mathcal{H}\left(-\bm{s}\right)=-\frac{1}{2}\left(-\bm{s}\right)^{\top}\bm{J}\left(-\bm{s}\right)=-\frac{1}{2}\bm{s}^{\top}\bm{J}\bm{s}=\mathcal{H}\left(\bm{s}\right)$. This symmetry of the probability distribution makes the mean state vector the zero vector, $\bm{s}\right>=\bm{0}$ because $\bm{s}\right>=\sum_{\bm{s}}p\left(\bm{s}\right)\bm{s}=\sum_{-\bm{s}}p\left(-\bm{s}\right)\left(-\bm{s}\right)=\sum_{\bm{s}}p\left(\bm{s}\right)\left(-\bm{s}\right)=-\left<\bm{s}\right>$. Thus, the covariance matrix becomes $\bm{\varSigma}=\sum_{\bm{s}}p\left(\bm{s}\right)\bm{s}\bm{s}^{\top}$. The off-diagonal element (covariance) is $\varSigma_{i,j}=\sum_{\bm{s}}p\left(\bm{s}\right)s_{i}s_{j}$, and the diagonal element (variance) is $\varSigma_{i,i}=\sum_{\bm{s}}p\left(\bm{s}\right)s_{i}^{2}=\sum_{\bm{s}}p\left(\bm{s}\right)=1$. Therefore, the correlation between spin $i$ and $j$ is the covariance between them, $\frac{\varSigma_{i,j}}{\sqrt{\varSigma_{i,i}}\sqrt{\varSigma_{j,j}}}=\varSigma_{i,j}$. The covariance matrix becomes the correlation matrix. PCA finds the most correlated direction in the hypercubic state space.
@@
Figure 10: Hypercubic energy landscape of finite artificial spin-ice systems and interaction networks colored by the angle of the biplot vectors. (a) The hypercubic energy landscape of one-ring (a1), two-ring (a2), and three-ring (a3) systems by PC1 and PC2. For clarity of the panel, only the states with the lowest and the second-lowest energy are shown. The energy of the state colors vertices (blue for the lowest energy and red for the second-lowest energy). Biplot vectors are shown on the top right, and their color indicates the angle of the biplot vectors. Notice that in (a2) and (a3), several biplot vectors are overlapped. One example pathway from a ground state to another ground state is shown as colored hypercubic edges in (a1). The colors of these lines are the same as the color of the corresponding biplot vectors. Examples of states are shown as insets in (a1), (a2), and (a3). See the caption of Fig. 9 for the definition of state visualization on the interaction network. (b) PC1 loading of one-ring (b1), two-ring (b2), and three-ring (b3) systems. The color of the bar matches the color of the corresponding biplot vectors in (a). (c) Same as (b) but for PC2. (d) The Ising spin interaction networks of one-ring (d1), two-ring (d2), and three-ring (d3) systems, where the nodes are colored by the angle of the corresponding biplot vectors in (a).
In Figs. 10(a1)–10(a3), the hypercubic energy landscapes of finite artificial kagome Ising spin systems are projected using PCA. Each state $\bm{s}$ corresponds to a vertex of the hypercube, and a state transition with a single spin flip corresponds to the edge of the hypercube. Thus, a hypercubic energy landscape is a kind of state transition diagram. For clarity of visualization, we show only the vertices with the lowest and the second-lowest energy, i.e., the ground states and the first-excited states, out of $2^{N}$ vertices (states) in Figs. 10(a1)–10(a3). By visualizing hypercubic edges connecting the ground states and first-excited states, one can observe the pathways expected to be followed in the dynamics at low temperatures. Note also that the distributions discussed in Sec. V appear here: the one-, two-, and three-ring systems have probability distributions that closely resemble the bipolar, quadripolar, and sexapolar cases, respectively.
A glance at the hypercubic energy landscapes in Figs. 10(a1)–10(a3) reveals a qualitative difference between the unfrustrated (one-ring) system and the frustrated (two- and three-ring) systems. In the one-ring system, the ground states are not directly connected by hypercubic edges; instead, they are always separated by excited states. In contrast, in the two- and three-ring systems, some ground states are directly connected by hypercubic edges. This shows that, in unfrustrated systems, ground states are always surrounded by excited states, whereas in frustrated systems, certain ground states are adjacent to each other. Thus, the degeneracy of ground states in frustrated systems qualitatively changes the structure of the hypercubic energy landscape.
VII.4 The hypercubic energy landscape of the one-ring system
@@
We first examine the one-ring system of Figs. 9(a1) and 9(b1). In Fig. 10(a1), we show the possible pathway from a ground state $\begin{bmatrix}\circ&\bullet&\bullet&\circ&\circ&\bullet\end{bmatrix}^{\top}$ to the globally spin-flipped state $\begin{bmatrix}\bullet&\circ&\circ&\bullet&\bullet&\circ\end{bmatrix}^{\top}$ by connecting the vertices with the lowest energy and the second-lowest energy using hypercubic edges. (See Figs. 9(b1) or 10(d1) for labeling of spins.) In this example, the hypercubic edge corresponds to the state transition with a single spin flip. The biplot vectors are shown at the top right of Fig. 10(a1), and its color indicates the angle. We emphasize one particular pathway between two ground states by coloring the edges following the biplot vectors in Fig. 10(b1). We found the state transition pathway is collective: the flipping order is restricted to keep the energy low. In a one-ring system, there are $6!=720$ possible pathways from one ground state to another ground state, but the number of most probable pathways is limited, as shown in Fig. 10(a1), and it is reduced to $966\times 2^{4}\times 1=$ 96 [52]. This reduction arises from the interaction network of the system and the associated hypercubic energy landscape.
@@
In addition, one can infer which spin is involved in which state transition by the biplot vectors. One can know which spin flips in the state transition by comparing the length and angle of the hypercubic edge of interest with that of the biplot vectors. The direction of the arrow of the biplot vector indicates the direction of the spin flip. In other words, if the state transition is along the direction of the arrow of the biplot vector, the spin flips from down to up. Otherwise, the spin flips from up to down. Notice that Fig. 10(a1) is the Hamming projection because the probability distribution is roughly the same as Eq. (13).
@@
To interpret the relation between PC loading and the interaction network, we visualize the loading of PC1 and PC2 in Figs. 10(b1) and 10(c1), respectively, colored by the angle of the biplot vector in Fig. 10(a1). We confirmed that the element-wise sign of PC1 corresponds to the ground state [Fig. 10(b1)]. The element-wise sign of PC2 loading corresponds to the state perpendicular to the ground state, which has a Hamming distance of $\frac{N}{2}$ from the ground state [Fig. 10(c1)].
@@
The interaction network in Fig. 10(d1), where nodes are colored by the angle of the biplot vectors, shows the correlation captured by the first two PC loadings. Spins 1 and 5 have the same color, and spin 3, interacting with both spins 1 and 5, has the color complementary to that of spins 1 and 5. We can argue the same for spins 2, 4, and 6. Hence, the resulting hypercubic energy landscape is drawn to emphasize the correlation arising from the interaction.
VII.5 The hypercubic energy landscape of the two-ring system
@@
The hypercubic energy landscape of the two-ring system, which has several ground states, is illustrated in Fig. 10(a2). We show six ground states and pathways between them in a two-ring system with a biplot vector on the top right. Similar to the one-ring system in Fig. 10(a1), the most probable pathways between the ground states are shown by connecting the vertices with the lowest and the second-lowest energy. We discuss two pathways connecting the ground states as examples.
@@
The second example is the transition pathway from the ground state on the upper left side (around $\begin{bmatrix}-3&0.3\end{bmatrix}^{\top}$), to the ground state on the top (around $\begin{bmatrix}0&3.2\end{bmatrix}^{\top}$), i.e., state transition from $\begin{bmatrix}\circ&\bullet&\circ&\bullet&\bullet&\bullet&\circ&\circ&\bullet&\circ&\bullet\end{bmatrix}^{\top}$ to $\begin{bmatrix}\bullet&\circ&\circ&\bullet&\circ&\bullet&\circ&\bullet&\circ&\circ&\bullet\end{bmatrix}^{\top}$, which requires the flipping of spins 1, 2, 5, 8, and 9 [see biplot vectors of Fig. 10(a2) and color of Figs. 10(b2), 10(c2), and 10(d2)]. This projection of the hypercube indicates that state transition happens with the correlated spin flips on the left-half part of the system (spins 1, 2, 5, 8, and 9). Among the left-half part of the system, spins 2 and 9 are more likely to flip at the beginning of the pathway. This difference arises from the unsatisfied interactions: $J_{3,9}$ and $J_{6,2}$. Spins 2 and 9 receive weaker local fields than spins 1, 5, and 8 because they have interactions involving spins 3, 6, and 10. Thus, spins 2 and 9 are expected to flip more than spins 1, 5, and 8. Because of the frustrated interactions involving spins 3, 6, and 10, spins 2 or 9 are the most probable to flip first. Then, the flips happen in the order of spins 2, 1, 5, 8, and 9 if spin 2 flips first, or in the order of spins 9, 8, 5, 1, and 2 if spin 9 flips first. Similar to the one-ring system in Fig. 10(a1), the number of most probable pathways is limited, but the constraints are stricter and fewer pathways are possible.
@@
To deepen our understanding of PC loading and interaction networks, we show the loading of PC1 and PC2 in Figs. 10(b2) and 10(c2), coloring them by the angles of the biplot vectors in Fig. 10(a2). Again, we confirmed that the sign of PC1 loading [Fig. 10(b2)] corresponds to the ground state with the largest magnitude of the PC1 score, but spin 6 is not determined. Thus, the frustrated dynamics of spin 6 are captured by PC1. The element-wise sign of PC2 loading [Fig. 10(c2)] corresponds to the ground state with the largest magnitude of the PC2 score. Notice that the sixth element of PC2 loading has a slightly higher magnitude than the others, indicating the importance of spin 6. If spin 6 flips when the system is in the ground state $\operatorname{sgn}\left(\bm{u}_{2}\right)$, the energy of the system increases more than when another spin flips.
@@
The interaction network with nodes colored by the angles of the biplot vectors in Fig. 10(d2) shows the inter-spin correlation captured by the biplot vectors. The biplot vectors capture the (anti)correlation of spins 1, 2, 5, 8, and 9 [yellow and blue arrows in Fig. 10(a2) top right] which corresponds to the left ring of the interaction network in Fig. 10(d2). The biplot vectors also capture the (anti)correlation of spins 3, 4, 7, 10, and 11 [green and purple arrows in Fig. 10(a2) top right] which corresponds to the right ring of the interaction network in Fig. 10(d2). The hypercubic energy landscape using PCA provides insight into the dynamics through the correlation of spins.
VII.6 The hypercubic energy landscape of the three-ring system